OpenStudy (anonymous):

rationalize the denominator of 3/square root of 2

OpenStudy (anonymous):

So when you are rationalizing, the easiest thing to do with equations like this is to multiply the numerator and the denominator by the square root number that is the denominator

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {3 \over \sqrt{2}} . {\sqrt{2} \over \sqrt{2}} }\)

OpenStudy (anonymous):

\[\frac{3}{\sqrt{2}} \times \frac{\sqrt{2}}{\sqrt{2}}\]

OpenStudy (anonymous):

there's always some1 faster :(

OpenStudy (anonymous):

that doesn't give the answer, that gives the concept of rationalizing

OpenStudy (anonymous):

basically the latex form of what you described, actually